# Forced 3D reconnection in an exponentially separating magnetic field

### v1

@@ Abstract
Matching to Eq. 11 requires $p_{0}=0$, $C_{0}=-\mathcal{J}_{0}/2\mathcal{S}$. Thus, the outer solution is
@@
Eqs. (9), (10), (19), and (20) are as obtained from numerical solution of Eqs. (5) and (6) (see Fig. 2). Qualitatively, the solution for $\tilde{t}\ll\tau_{\mathrm{rec}}$ [the $n=0$ terms in Eqs. (9) and (10)] is a self-similar spreading with similarity variable $\xi$. The solution is diffusive, but corresponds to negligible reconnection [see Eq. (22) below] until the current begins to decay at $\tilde{t}\sim\tau_{\mathrm{rec}}$, which is when the propagating front reaches ~$\tilde{\ell}\sim\ln\mathcal{S}$ (corresponding to $\xi\sim 1$).
@@
MHD case: (v) Matching solutions, $\tilde{t}\gtrsim\mathcal{S}$. If ${\tilde{t}\gtrsim\mathcal{S}}$, we cannot neglect $\tilde{t}$ in the logarithm in Eq. (17), nor the final term in Eq. (10). We show how to modify the solution in Section 4 of the Supplementary Information  [^3] Qualitatively, the effect of the modification is that decay of $\tilde{u}_{\mathrm{out}}$ and $\tilde{J}_{\parallel\mathrm{out}}$ is arrested at a non-zero value.. However, because $\tau_{\mathrm{rec}}\ll\mathcal{S}$ as $\mathcal{S}\to\infty$, the solution in the preceding sections is the one of primary interest.
Kinetic case: (i) Inner solution (ideal). We neglect inertia in the inner region—because this is the only term in which the kinetic equations are different from the MHD ones, the solutions Eqs. (9) and (10) remain valid in the kinetic case [although $\tilde{J}_{\parallel 0}(\tilde{t})$ is different].
Kinetic case: (ii) Outer solution (diffusive). In the kinetic case, the governing equation for $\tilde{u}_{\mathrm{out}}$ is a diffusion equation with constant coefficient $\mathcal{S}$. We recast it in a form similar to Eq. (23) by changing variables from $\tilde{\ell}$ to a new similarity variable ${\xi=\tilde{\ell}/(4\mathcal{S}\tilde{t})^{1/2}}$:
$\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{t}}=\mathcal{S}\frac{\partial^{2}\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}^{2}}\,\,\to\,\,\tilde{t}\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{t}}=\frac{1}{4}\frac{\partial^{2}\tilde{u}_{\mathrm{out}}}{\partial\xi^{2}}+\frac{1}{2}\xi\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\xi}.$ (23)
The solutions of the ordinary differential equation that results from substituting Eq. (14) into Eq. (23) are [cf. (15)] ${f_{n}(\xi)=C_{n}e^{-\xi^{2}}U\left((1+2p_{n})/2,1/2,\xi^{2}\right)}$.
Figure 3: Evolution of $\Delta\psi$ [Eq. (8)] as a function of time in the MHD (left) and kinetic (right) cases in the numerical solution of Eqs. (5) and (6). At $\tilde{t}\gg 1/\epsilon$, the evolution is as predicted by Eqs. (22) and (28) (see insets, which show time rescaled by the appropriate function of $\mathcal{S}$). Blue regions indicate $\tilde{t}\lesssim\epsilon^{-1}$, when the force $\tilde{F}$ has not yet reached its final value. The power laws in these regions are derived in Section 3 of the Supplementary Information.
Kinetic case: (iii) Matching solutions. The matching condition is Eq. (16), as in the MHD case. From the small-argument limit of Tricomi’s function, we have
where
These solutions are as we find numerically (see Fig. 2). Qualitatively, the solution for $\tilde{t}\ll\tau_{\mathrm{rec}}$ is a self-similar spreading of both $\tilde{u}_{\mathrm{out}}$ and $\tilde{J}_{\parallel\mathrm{out}}$, with $\tilde{u}_{\mathrm{out}}$ also growing like $\tilde{t}^{1/2}$. Reconnection occurs when the current begins to decay at $\tilde{t}\sim\tau_{\mathrm{rec}}$ [Eq. (28)], which is when the propagating front reaches ~${\tilde{\ell}\sim\mathcal{S}^{2/3}}$.
Kinetic case: (iv) Reconnection rate. Taking the ${\xi\to 0}$ limit of Eq. (26), we find [cf. Eq. (22)]
@@
where $E_{3/2}(x)$ is the Mittag–Leffler function, a generalization of the exponential function defined by ${E_{\alpha}(x)=\sum_{k=0}^{\infty}x^{k}/\Gamma(1+\alpha k)}$. Eq. (28) reproduces the numerical evolution of $\Delta\psi$, shown in Fig. 3.
@@
Generalization to other geometries. In the kinetic case, the outer-region diffusion equation (23) is independent of |~$|\tilde{\bm{\nabla}}\alpha|(\ell)$. It follows that the solution (28) is valid for any magnetic geometry, not just the one described by Eq. (2). By contrast, |~$|\tilde{\bm{\nabla}}\alpha|$ does appear in Eq. (13) for the MHD case. For an equilibrium with |~$|\tilde{\bm{\nabla}}\alpha|\propto\tilde{\ell}^{\delta}$ for ~$\tilde{\ell}\gg 1$, it is readily verified that Eq. (13) has a similarity variable $\xi=\tilde{\ell}/(\mathcal{S}\tilde{t})^{1/2(\delta+1)}$. It follows that  [^4] Here we have used the relations $\tilde{J}_{\parallel 0}=\mathcal{S}\lim_{\tilde{\ell}\to 0}\partial\tilde{u}_{\mathrm{out}}/\partial\tilde{\ell}$, $\mathrm{d}\tilde{J}_{\parallel 0}/\mathrm{d}\tilde{t}=\lim_{\tilde{\ell}\to 0}\tilde{u}_{\mathrm{out}}$ and $\lim_{\tilde{t}\to 0}\tilde{u}_{\mathrm{out}}\propto\tilde{t}^{p_{0}}f_{0}(\xi)$ [Eq. (14)], which imply that $\mathrm{d}\ln J_{\parallel 0}/\mathrm{d}\tilde{t}\propto\mathcal{S}^{1/2(\delta-1)}\tilde{t}^{1/2(\delta+1)}$ as $\tilde{t}\to 0$
@@
$\Delta\psi\to\left(\frac{\tilde{t}}{\tau_{\mathrm{rec}}}\right)^{(\delta+3/2)/(\delta+1)},\quad\tau_{\mathrm{rec}}\propto\mathcal{S}^{(2\delta+1)/(2\delta+3)}$ (29)
as $\tilde{t}/\tau_{\mathrm{rec}}\to 0$. Eq. (29) implies that MHD reconnection becomes slower as the rate of field-line separation increases (i.e., as $\delta$ increases). In the case of linear shear, i.e., $\delta=1$, for example, $\tau_{\mathrm{rec}}\propto\mathcal{S}^{3/5}$ (as was inevitable because Eqs. (5) and (6) are then isomorphic to the 2D tearing-mode equations [11]). We note that the time taken for the diffusive solutions to develop increases as $\delta\to 0$; this point will be discussed in a future publication.
@@
Conclusion.  We have derived nonlinear solutions for 3D MHD and electron-only reconnection in a sheared magnetic field. In our solutions, reconnection is induced by diffusing Alfvén- and semi-collisional kinetic-Alfvén-wave packets generated by an external force applied gradually to exponentially separating field lines. Their diffusion changes the connectivity of the global field in a timescale $\propto\mathcal{S}/\ln\mathcal{S}$ and $\propto\mathcal{S}^{1/3}$, respectively. Interestingly, the former timescale reduces if field lines separate more slowly than exponentially with distance [Eq. (29)], while the latter is unchanged.
@@
Acknowledgements. We are grateful to Christopher Ham and Anna Thackray for their collaboration on a related project. We also thank the hosts, organisers and participants of the 2024 Plasma Kinetics Workshop at the Wolfgang Pauli Institute, Vienna, where this work was discussed.
Supplementary information
@@
where $L=L(\alpha,\psi)$ is arbitrary and
@@
.2 2. Derivation of the dynamical equations
@@
.3 3. Evolution of $\Delta\psi$ at $\epsilon\tilde{t}\sim 1$.
At early times $\epsilon\tilde{t}\sim 1$, Eq. (16) is revised to
@@
Substituting Eq. (9) into Eq. (8), we have that
@@
where we have taken $\tilde{F}=(1-e^{-\epsilon^{2}\tilde{t}^{2}})\tilde{f}(\ell)$, for some function $\tilde{f}(\ell)$ ($F_{0}\tilde{l}^{3}e^{-\ell^{2}}$ in the plots shown in the main text).
@@
.4 4. MHD solution for $\tilde{t}\sim\mathcal{S}$
Including the resistive term in (10), Eq. (16) becomes (at $\epsilon^{-1}\ll\tilde{t}$) [using Eq. (9) and ${\tilde{J}_{\parallel\mathrm{out}}=\mathcal{S}\partial\tilde{u}_{\mathrm{out}}/\partial\tilde{\ell}}$],
$\lim_{\tilde{\ell}\to 0}\tilde{u}_{\mathrm{out}}=\frac{\mathcal{S}}{|\Delta^{\prime}|}\frac{\mathrm{d}}{\mathrm{d}\tilde{t}}\lim_{\tilde{\ell}\to 0}\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}}+\tilde{u}_{F\mathcal{S}}+\zeta(\tilde{\ell})\lim_{\tilde{\ell}\to 0}\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}}$ (54)
Substituting Eq. (17) into Eq. (54) yields

### v2

@@ Abstract
Finally, Eq. (11) requires $p_{0}=0$, $C_{0}=-\mathcal{J}_{0}/2\mathcal{S}$, whence
@@
Eqs. (9), (10), (19), and (20) are as we obtain from numerical solution of Eqs. (5) and (6) (see Fig. 2). Qualitatively, the solution for $\tilde{t}\ll\tau_{\mathrm{rec}}$ [the $n=0$ terms in Eqs. (19) and (20)] is a self-similar spreading with similarity variable $\xi$. The solution is diffusive, but corresponds to negligible reconnection [Eq. (22)] until the current starts to decay at $\tilde{t}\sim\tau_{\mathrm{rec}}$, which is when the propagating front reaches ~$\tilde{\ell}\sim\ln\mathcal{S}$ (corresponding to $\xi\sim 1$).
@@
MHD case: (v) Matching solutions, $\tilde{t}\gtrsim\mathcal{S}$. Combining Eqs. (10) and (21), we find that the contribution of the final term (involving $\tilde{F}$) in (10) to $c_{1}(\tilde{t})$ is negligible only for $\tilde{t}\ll\mathcal{S}$—at later times, this term must also be included in the matching condition (16). Physically, at $\tilde{t}\sim\mathcal{S}$, resistivity can no longer be neglected at the location of the force, and the flux tube begins to “slip” diffusively there. We derive the corresponding solution in Section 4 of the End Matter  [^3] Qualitatively, the effect of the modification is that decay of $\tilde{u}_{\mathrm{out}}$ and $\tilde{J}_{\parallel\mathrm{out}}$ is arrested at a non-zero value.. However, because $\tau_{\mathrm{rec}}\ll\mathcal{S}$ as $\mathcal{S}\to\infty$, the solution in the preceding sections is the one of primary interest.
Kinetic case: The derivation of the reconnection rate in the kinetic case of Eq. (5) is analogous to the MHD case; we present it in Section 2 of the End Matter. The chief difference is that, unlike Eq. (12), the diffusion equation for the outer region in the kinetic case [Eq. (32)] is independent of $|\bm{\nabla}\alpha|$, so the diffusion coefficient is not suppressed by field-line separation. We find [cf. Eq. (22)]
@@
where $\tau_{\mathrm{rec}}=(\mathcal{S}/|\Delta^{\prime}|^{2})^{1/3}$ and $E_{3/2}(x)$ is the Mittag–Leffler function, a generalization of the exponential function defined by ${E_{\alpha}(x)=\sum_{k=0}^{\infty}x^{k}/\Gamma(1+\alpha k)}$.
@@
Generalization to other geometries. As explained above, the outer-region diffusion equation is independent of |~$|\tilde{\bm{\nabla}}\alpha|(\ell)$ in the kinetic case. The solution (23) is therefore valid for any magnetic geometry, not just the one described by Eq. (2). By contrast, |~$|\tilde{\bm{\nabla}}\alpha|$ does appear in Eq. (13) (MHD case). For an equilibrium with |~$|\tilde{\bm{\nabla}}\alpha|\propto\tilde{\ell}^{\delta}$ for ~$\tilde{\ell}\gg 1$, it is readily verified that Eq. (13) has a similarity variable $\xi=\tilde{\ell}/(\mathcal{S}\tilde{t})^{1/2(\delta+1)}$. It follows that  [^4] In deriving these expressions, we have used the relations ${\tilde{J}_{\parallel 0}=\mathcal{S}\lim_{\tilde{\ell}\to 0}\partial\tilde{u}_{\mathrm{out}}/\partial\tilde{\ell}}$, ${\mathrm{d}\tilde{J}_{\parallel 0}/\mathrm{d}\tilde{t}=\lim_{\tilde{\ell}\to 0}\tilde{u}_{\mathrm{out}}}$ and ${\lim_{\tilde{t}\to 0}\tilde{u}_{\mathrm{out}}\propto\tilde{t}^{p_{0}}f_{0}(\xi)}$ [Eq. (14)], which together imply that ${\mathrm{d}\ln J_{\parallel 0}/\mathrm{d}\tilde{t}\propto{\color[rgb]{0,0,0}\mathcal{S}^{-(1+2\delta)/2(1+\delta)}}\tilde{t}^{1/2(\delta+1)}}$ as $\tilde{t}\to 0$
@@
$\Delta\psi\to\left(\frac{\tilde{t}}{\tau_{\mathrm{rec}}}\right)^{\tfrac{\delta+3/2}{\delta+1}},\quad\tau_{\mathrm{rec}}\propto\frac{\mathcal{S}^{(2\delta+1)/(2\delta+3)}}{|\Delta^{\prime}|^{2(\delta+1)/(2\delta+3)}}$ (24)
as $\tilde{t}/\tau_{\mathrm{rec}}\to 0$. Eq. (24) implies that MHD reconnection becomes slower as the rate of field-line separation increases (i.e., as $\delta$ increases). In the case of linear shear, i.e., $\delta=1$ (as for field lines not on the separatrix) $\tau_{\mathrm{rec}}\propto\mathcal{S}^{3/5}$  [^5] This dependence was inevitable as, in the case of linear shear, Eqs. (5) and (6) are isomorphic to the 2D tearing-mode equations [24].. Faster reconnection occurs for field lines that separate more slowly (decreasing $\delta$), but only up to $\delta=1/2$, as then $|\Delta^{\prime}|$ diverges [see its definition below (11)]. Physically, $\delta=1/2$ is the smallest value of $\delta$ for which the diffusive solution can be causally connected to the origin at $t\sim\tau_{\mathrm{rec}}$. The shortest possible MHD-reconnection timescale in our theory as $\mathcal{S}\to\infty$ is therefore $\tau_{\mathrm{rec}}\sim\mathcal{S}^{1/2}/|\Delta|^{3/4}$, corresponding to $\delta\to 1/2^{+}$.
@@
Conclusion.  We have derived nonlinear solutions for 3D MHD and electron-only reconnection in a sheared magnetic field. In our solutions, reconnection is induced by diffusing Alfvén- and semi-collisional kinetic-Alfvén-wave packets generated by an external force applied gradually to exponentially separating field lines. Their diffusion changes the connectivity of the global field in timescales $\propto\mathcal{S}/\ln\mathcal{S}$ and $\propto\mathcal{S}^{1/3}$, respectively. We expect our results to be quantitatively applicable to reconnection at separatrices, as in tokamak-edge-field schochasticization by an ELM-control (RMP) coil, although we note that the anisotropic ordering $k_{\parallel}\gg k_{\perp}$ necessary for a tractable theory is not rigorously applicable in that case. Qualitatively, our paper provides a solvable case of 3D magnetic reconnection, with the interesting and unexpected property that the reconnection rate is greater for field lines that diverge more slowly [Eq. (24)] for MHD dynamics, and is geometry-independent in the case of semi-collisional electron-only reconnection [21].
@@
Acknowledgements. We are grateful to Christopher Ham and Anna Thackray for their collaboration on a related project. We also thank the hosts, organizers, and participants of the 2024 Plasma Kinetics Workshop at the Wolfgang Pauli Institute, Vienna, where this work was discussed.
End matter
@@
where the lower limit $L=L(\alpha,\psi)$ is arbitrary and
@@
.2 2. Solution for the kinetic case
Kinetic case: (i) Inner region (~$(\tilde{\ell}\ll\mathcal{S}^{2/3})$. We neglect inertia in the inner region (as justified a posteriori in Section 3 of the End Matter). This is the only term in which the kinetic and MHD cases of Eq. (5) differ, so Eqs. (9) and (10) remain valid for the kinetic case.
Kinetic case: (ii) Outer region (~$(\tilde{\ell}\gg\ln\mathcal{S})$.  As justified a posteriori in Section 3 of the End Matter, we now neglect the forcing term in Eq. (5) and the induction term in Eq. (6). Then, $\tilde{u}_{\mathrm{out}}$ obeys the diffusion equation [cf. Eq. (12) for the MHD case]
$\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{t}}={\mathcal{S}}\frac{\partial^{2}\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}^{2}}.$ (32)
We recast Eq. (32) in a form similar to Eq. (13) by changing variables from $\tilde{\ell}$ to a new similarity variable ${\xi=\tilde{\ell}/(4\mathcal{S}\tilde{t})^{1/2}}$, which yields
$\tilde{t}\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{t}}=\frac{1}{4}\frac{\partial^{2}\tilde{u}_{\mathrm{out}}}{\partial\xi^{2}}+\frac{1}{2}\xi\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\xi}.$ (33)
The solutions of the ordinary differential equations that result from substituting Eq. (14) into Eq. (33) are ${f_{n}(\xi)=C_{n}e^{-\xi^{2}}U\left((1+2p_{n})/2,1/2,\xi^{2}\right)}$.
Kinetic case: (iii) Matching solutions. Because Eq. (6) is the same in the MHD and kinetic cases, Eq. (16) is the matching condition for both [see below Eq. (38) for justification of neglect of the second two terms in Eq. (10) in forming the matching condition]. From the small-argument limit of Tricomi’s function, we have
where
These solutions are as we find numerically (Fig. 4). Qualitatively, both $\tilde{u}_{\mathrm{out}}$ and $\tilde{J}_{\parallel\mathrm{out}}$ spread self-similarly for ${\tilde{t}\ll\tau_{\mathrm{rec}}}$, with $\tilde{u}_{\mathrm{out}}$ also growing like $\tilde{t}^{1/2}$. Reconnection occurs when the current begins to decay at $\tilde{t}\sim\tau_{\mathrm{rec}}$, which is when the propagating front reaches ~${\tilde{\ell}\sim\mathcal{S}^{2/3}}$. The ${\xi\to 0}$ limit of Eq. (36) yields
$\tilde{J}_{\parallel 0}=\mathcal{J}_{0}E_{3/2}\left[-\left(\tilde{t}/\tau_{\mathrm{rec}}\right)^{3/2}\right],$ (38)
where $E_{3/2}(x)$ is the Mittag–Leffler function. We note that $\tilde{J}_{\parallel 0}=O[(\tilde{t}/\tau_{\mathrm{rec}})^{-3/2}]$ as $\tilde{t}/\tau_{\mathrm{rec}}\to\infty$, from which it follows that the term involving $\tilde{F}$ in Eq. (10) does not contribute to $c_{1}$ [see “MHD case: (iii)”] until ${\tilde{t}\sim\mathcal{S}\gg\tau_{\mathrm{rec}}}$, and therefore is always negligible in forming the matching condition (16) [cf. “MHD case: (iv)”]. Combining Eqs. (8) and (38) yields Eq. (23) for $\Delta\psi$.
Figure 4: As Fig. 2, but for the kinetic case.
.3 3. A posteriori justification of neglected terms in matched asymptotics
MHD case: Inner region. From Eqs. (9) and (10), we have that, for $\epsilon\tilde{t}\gg 1$ and ~$\tilde{\ell}\gg 1$,
$\frac{\partial\tilde{u}_{\mathrm{in}}}{\partial\tilde{t}}=\frac{1}{B_{0}}\int_{0}^{\ell}\mathrm{d}\ell^{\prime}\frac{B_{0}}{|\bm{\nabla}\alpha|^{2}}\left(\frac{\mathrm{d}^{2}\tilde{J}_{\parallel 0}}{\mathrm{d}\tilde{t}^{2}}+\frac{|\bm{\nabla}\alpha|^{2}}{\mathcal{S}}\frac{\mathrm{d}\tilde{J}_{\parallel 0}}{\mathrm{d}\tilde{t}}{}\right).$ (39)
Using Eq. (21) for $\tilde{J}_{\parallel 0}$ in Eq. (39), the neglected term involving $\partial\tilde{u}_{\mathrm{in}}/\partial\tilde{t}$ in Eq. (9) is small compared to $\tilde{J}_{\parallel 0}$ if
~|~$\int_{-\infty}^{\tilde{\ell}}\frac{|\tilde{\bm{\nabla}}\alpha|^{2}}{B_{0}}\left(\frac{1}{|\Delta^{\prime}|\tau_{\mathrm{rec}}}-\frac{\tilde{\ell}}{\mathcal{S}}\right)\frac{1}{\tau_{\mathrm{rec}}}\ll 1.$ (40)
This is true for ~$\tilde{\ell}\ll\ln\mathcal{S}$, which is therefore the domain of validity of the inner solution.
MHD case: Outer region. The induction term in Eq. (6) is negligible compared to the diffusive one for sufficiently small $\tilde{\ell}$ because of its $1/|\tilde{\bm{\nabla}}\alpha|^{2}$ prefactor [from Eq. (20) (visualized in Fig. 2), $\partial\ln\tilde{J}_{\parallel\mathrm{out}}/\partial t$ has no strong dependence on $\tilde{\ell}$ to prevent this]. Assuming that the outer solution is valid for ~$\tilde{\ell}\ll\ln\mathcal{S}$, then matching to the inner condition yields $\tilde{J}_{\parallel\mathrm{out}}=\tilde{J}_{\parallel 0}$ there. It follows that the ratio of the induction to resistive terms, $({\mathcal{S}/|\tilde{\bm{\nabla}}\alpha|^{2}})\partial\ln\tilde{J}_{\parallel\mathrm{out}}/\partial t$, is $\ln\mathcal{S}/|\tilde{\bm{\nabla}}\alpha|^{2}$ at ~$\tilde{\ell}\ll\ln\mathcal{S}$. The induction term is therefore negligible for ~$\tilde{\ell}\gg\ln(\ln\mathcal{S})$, which is the domain of validity of the outer solution (this includes ~$\tilde{\ell}\ll\ln\mathcal{S}$, as assumed).
Kinetic case: Inner region. Eq. (39) holds for the kinetic case; substituting Eq. (38) and using Eq. (37) for $\tau_{\mathrm{rec}}$, it follows from an analogous condition to (40) that $\partial\tilde{u}_{\mathrm{in}}/\partial\tilde{t}$ is negligible in Eq. (9) for ~${\tilde{\ell}\ll\mathcal{S}^{2/3}}$. This is the domain of validity of the inner solution.
Kinetic case: Outer region. The ratio of induction and diffusion terms in (6) is $\mathcal{S}/(\tau_{\mathrm{rec}}|\bm{\nabla}\alpha|^{2})$. This is small provided that $\ell\gg\ln\mathrm{S}$, which is therefore the domain of validity of the outer solution.
.4 4. MHD solution for $\tilde{t}\sim\mathcal{S}$
Retaining all terms in Eq. (10), Eq. (16) becomes, for $\epsilon\tilde{t}\gg 1$ [using Eq. (9) and ${\tilde{J}_{\parallel\mathrm{out}}=\mathcal{S}\partial\tilde{u}_{\mathrm{out}}/\partial\tilde{\ell}}$],
$\lim_{\tilde{\ell}\to 0}\tilde{u}_{\mathrm{out}}=\frac{\mathcal{S}}{|\Delta^{\prime}|}\frac{\mathrm{d}}{\mathrm{d}\tilde{t}}\lim_{\tilde{\ell}\to 0}\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}}+\tilde{u}_{F\mathcal{S}}+\zeta(\tilde{\ell})\lim_{\tilde{\ell}\to 0}\frac{\partial\tilde{u}_{\mathrm{out}}}{\partial\tilde{\ell}},$ (41)
Substituting Eq. (17) into Eq. (41) yields
Supplementary information
.5 1. Derivation of the dynamical equations
@@
.6 2. Evolution of $\Delta\psi$ at $\epsilon\tilde{t}\sim 1$.
For $\epsilon\tilde{t}\sim 1$ (and $\tilde{t}\ll\mathcal{S}$), Eq. (10) has an additional term involving the time derivative of $\tilde{F}$. Including this term, Eq. (16) becomes
@@
We now describe how the inclusion of this term changes evolution of $\Delta\psi$, which, from substituting Eq. (9) into Eq. (8), is given by
@@
where we have taken $\tilde{F}=(1-e^{-\epsilon^{2}\tilde{t}^{2}})\tilde{f}(\ell)$, for some function $\tilde{f}(\ell)$ (chosen to be $\propto\tilde{l}^{3}e^{-\ell^{2}}$ in the plots shown in the main text).